88456:

An unopened cereal box contains `15/28` cubic foot of cereal. If the box is full, and the length and height of the box are `5/7` and `1(1)/(2)` feet, respectively, what is the width of the box? A. `1/2` foot B. `1/7` foot C. `3/5` foot D. `4/7` foot

88456:

pls help, idk how 2 do it.

angiesyl:

15/28=(5/7)(1 1/2)(w) 15/28=(5/7)(3/2)w 15/28=(15/14)w (15/28)/(15/14)=w (15/28)*(14/15)=w w=14/28 w=1/2 ft

angiesyl:

hope tha helped